# SSP-GNN: Learning to Track via Bilevel Optimization thanks: Corresponding author

### v2

@@ Abstract
Index Terms:
feature-aided tracking, data association, GNN
@@ I Introduction
Learning-based approaches to tracking [31, 26] learn key components of a tracker (e.g., data association) directly from annotated training data. Recently, approaches to neural message passing [11] have been adopted by [5] to formulate tracking on a detection graph (where a detection graph is instantiated for a temporal window given all the measurements/detections). However, predicting which pairs of detections should be linked without inferring complete tracks, while computationally efficient, may be suboptimal.
We propose an end-to-end bilevel formulation for learning MOT, where the inner optimization (tracking) is accomplished via a successive shortest paths (SSP) algorithm on a tracking graph [6, 32] with edge costs. Given these costs, SSP finds a solution that is guaranteed to be optimal. In our outer optimization problem we learn parameters of a function that computes the cost of each edge in the tracking graph given detections that include attributes. The edge-cost prediction function takes the form of a graph neural network (GNN).
Learning GNN parameters requires solving a bilevel optimization problem that is different from those proposed earlier, e.g. [26, 16]. Specifically, we derive a novel loss function defined with respect to SSP-computed tracks and the ground-truth tracks and an algorithm to learn the GNN parameters by gradient descent. At each iteration, GNN parameters are updated to increase the cost of incorrect tracks, effectively learning from (tracking) mistakes.
@@
  • A novel end-to-end learnable approach to graph-based tracking with lower computational cost than [26, 16].
@@ II Related Work
Learning approaches to tracking In many domains, using machine-learning tools to find optimal parameters for a tracking algorithm from data have been shown to outperform hand-crafting and hand-tuning an algorithm. In the computer vision domain, learning-based approaches to multi-object tracking have taken many different forms. [31] Demonstrated a Markov decision process (MDP) capable of initiating and terminating target tracks; parameters of this MDP were learned from data. In [20] recurrent neural networks were used in an end-to-end framework. Specifically, in [20] a target’s state was maintained by a dedicated recurrent neural network (RNN); assignment of measurements to target tracks and track creation was estimated via a long-short-term memory (LSTM) neural network. A transformer neural network was used in [19] as a unified formulation for detecting (new) targets, extending tracks, and creating new tracks.
@@
Graph-based tracking via classical methods Application of graph-based and linear-programming (LP) methods to multi-target tracking has a long history, e.g., [6], and includes recent attempts to combine graph-based tracking with multi-hypothesis tracking (MHT) [9]. Graph-based tracking has shown to be effective in computer vision where it is often employed with a tracking-by-detection paradigm where tracking is reduced to the grouping of detections (bounding boxes) proposed by a (target class specific) object detector.
@@
An early graph-based method of [32] proposed the general form of a tracking graph and solved for an unknown number of targets via repeated invocation of min-cost flow. The computational complexity was reduced in [24] by introducing augmenting paths, thus extending an existing tracking solution to include an additional target instead of starting from scratch. While [32] and [24] performed tracking in the image plane, tracking on the ground plane from multiple camera views via K-shortest-paths (KSP) algorithm was proposed in [3]. Approaches mentioned thus far assumed that targets dynamics are Markovian and targets move independently of each other. An approach to extend LP-based tracking to take into account long-term (group) behavior was proposed in [18]. During training, their method alternated between learning behavior patterns and fitting trajectories to data given the observations and priors given by the learned patterns.
@@
Graph-based tracking via neural methods A significant limitation of early graph-based approaches [32, 24, 3] is the requirement to hand-craft scalar costs on tracking graph edges. However, in many real-world scenarios where we wish to apply feature-aided tracking, measurements (detections) in addition to containing positional information also include (high-dimensional) vectors of attributes, such as re-identification (ReID) features. Neural networks could learn a mapping from these high-dimensional attributes to scalar edge weights, but deriving an end-to-end learnable method is nontrivial. Among the first attempts to apply neural methods to graph-based tracking, the approach of [26] stands out for formulating learning and tracking as bilevel [33] optimization. Specifically, in [26] the high-level optimization problem solved for the neural network parameters, while the lower-level problem was a constrained LP that solved a tracking-by-detection problem over a temporal window of pre-defined length. The bilevel approach of [26] was revisited in [16], but instead of relying on the implicit function theorem to deal with the inner LP problem, the latter used KKT conditions to compute the gradient at the lower problem’s optimum.
@@
Recent advances in neural methods for graph-structured problems, such as graph neural networks (GNNs) [10] led to advances in graph-based tracking. Inspired by the neural message-passing on graphs method of [11], the authors of [5] proposed to learn a function to estimate edge probabilities in a tracking graph. This approach was extended in [7] to jointly solve for short-term and long-term tracking. While the approaches of [5, 7] require a batch of frames, and thus operate with a constant lag, approaches of [30, 19] require only the current frame to manage tracks. The ideas of [11] were adopted in [17] to develop a neural-enhanced message-passing data association algorithm.
@@
Neural approaches to optimization Learning-based approaches to MOT mentioned thus far solve the underlying assignment problem using classical methods e.g., LP. An approach to solve quadratic optimization as a layer in a neural network was proposed in [2]. In [23], the matrix of association costs was updated using Sinkhorn-Knopp’s algorithm [27] to satisfy MOT constraints. A differentiable graph matching layer was proposed in [12] and used to associate graphs formed on detections across time frames.
@@
Our contribution We draw on [10, 11] and [5] to formulate a GNN that infers edge costs on a graph structure, but unlike [5] our inner optimization with SSP ensures that track predictions are globally optimal and satisfy constraints. We demonstrate in Section IV that our method compares favorably to a GNN-based method without such inner optimization. Our method is similar to that of [16] in forming a bilevel problem during training, but we show in Section III-D that our inner optimization with SSP has lower computational complexity.
@@ III-A Graph-based formulation
Within the batch of $T$ data frames, we solve the tracking problem by casting it in a graph-based formulation [32, 24]. Given a temporal window $t\in 1,\ldots,T$ our objective is to group $\{{\bf y}_{j,t}\}$ into a set $\mathcal{P}\doteq\{p_{k}\}$ of paths in a tracking graph $\mathcal{G}^{\textrm{trk}}$ with positive and negative edge costs. The set $\mathcal{P}$ may be empty if the tracker doesn’t predict any tracks.
@@ III-B The detection graph and the tracking graph
Given detections $\mathcal{Y}\doteq\{{\bf y}_{j,t}\}_{t=1}^{T}$ in a temporal window with $T$ frames, we construct a detection graph $\mathcal{G}^{\textrm{det}}$. Nodes in $\mathcal{G}^{\textrm{det}}$ are in one-to-one correspondence with detections in $\mathcal{Y}$. An edge $(j,j^{\prime})$ exists between ${\bf y}_{j,t}$ and ${\bf y}_{j^{\prime},t^{\prime}}$ if $t<t^{\prime}\leq t+\Delta$ and the location components of ${\bf y}_{j,t}$ and ${\bf y}_{j^{\prime},t^{\prime}}$ satisfy gating constraints. The use of $\Delta>1$ is standard [26, 16] to handle short-duration obscuration (occlusion) of a target. In Fig. 2(a) we show an example of $\mathcal{G}^{\textrm{det}}$ for $T=3$ and $\Delta=2$. Two paths that satisfy node-disjoint constraints are shown.
Given $\mathcal{G}^{\textrm{det}}$, we define its companion tracking graph $\mathcal{G}^{\textrm{trk}}$, whose topology follows [6, 32]. In particular, every node in $\mathcal{G}^{\textrm{det}}$ is represented by a pair of “twin” nodes in $\mathcal{G}^{\textrm{trk}}$. Furthermore, $\mathcal{G}^{\textrm{trk}}$ contains source ${\bf s}$ and terminal ${\bf t}$ nodes with edges from ${\bf s}$ and to ${\bf t}$ defined in a way as to allow every detection to either correspond to a tracks’s birth or its termination. Fig. 2(b) shows the tracking graph $\mathcal{G}^{\textrm{trk}}$ derived from its $\mathcal{G}^{\textrm{det}}$; the paths (tracks) in $\mathcal{G}^{\textrm{trk}}$ originate in ${\bf s}$, pass through “twin” nodes and terminate in ${\bf t}$.
@@
Our $f_{\boldsymbol{\theta}}$ takes the form of a message-passing network (MPN) [11], a specialization of a graph neural net (GNN) [10]. We adopt the ideas of [11] and  [14] as implemented in [5].
@@
1 Initialize ${\bf h}^{(0)}_{(i,j)}$ and ${\bf h}^{(0)}_{z}$ for all $(i,j)\in\mathcal{E},z\in\mathcal{V}$
@@ 18       $c_{(i,j)}=g_{e}^{\mathrm{readout}}({\bf h}^{(L)}_{(i,j)},{\bf h}^{(L)}_{(j,i)})$
Messages are computed using the functions $g_{e}(\cdot,\cdot,\cdot)$ and $g_{v}(\cdot,\cdot)$ which take the form of learnable multi-layer perceptions. In our implementation, as suggested by [5], $g_{v}(\cdot,\cdot)$ consists of two independently parameterized functions $g_{v}^{\mathrm{past}}(\cdot,\cdot),g_{v}^{\mathrm{fut}}(\cdot,\cdot)$ which are specialized for forward and backward in time messages respectively. The readout function $g_{e}^{\mathrm{readout}}(\cdot,\cdot)$, also a learnable MLP, then computes the edge cost by aggregating the final attributes of the forward and backward in time edges between each pair of nodes. The parameters of these four functions comprise the vector ${\boldsymbol{\theta}}$.
@@ III-C End-to-end learning for tracking
Bilevel optimization to learn MPN parameters We formulate the following bilevel optimization problem [8, 33, 21] to find the optimal set of GNN parameters given ground-truth paths $\mathcal{P}^{+}$ :
@@
$\dfrac{\partial\mathcal{L}^{\textrm{SSP-GNN}}}{\partial{\boldsymbol{\theta}}}=\dfrac{\partial\ell}{\partial{\bf c}}\Big{|}_{\mathcal{P}^{*}}\dfrac{\partial{\bf c}}{\partial{\boldsymbol{\theta}}}\;,$ (4)
@@
In practice, we implement optimization of ${\boldsymbol{\theta}}$ in Equation 3 via a two-stage learning algorithm. In Stage I, inspired by [22] which introduced MCMC data association moves, we “perturb” $\mathcal{P}^{+}$ to pre-compute a fixed-size set of negative examples $\mathcal{P}^{-}$. Generating $\mathcal{P}^{-}$ does not require running the SSP algorithm.
@@ 9 ${\boldsymbol{\theta}}\leftarrow\mathsf{init\_gnn\_weights}()$
11       $\{c_{e}\}_{e=1}^{|\mathcal{E}|}\leftarrow f_{\boldsymbol{\theta}}(\mathcal{G})$
@@
17       $\{c_{e}\}_{e=1}^{|\mathcal{E}|}\leftarrow f_{\boldsymbol{\theta}}(\mathcal{G})$
@@
III-D Computational Complexity
In our analysis, to provide a fair comparison between our proposed algorithm and related methods we focus only on forward inference of the inner optimization problem used to compute $\mathcal{P}^{\star}$ from $\mathcal{G}^{\mathrm{trk}}$. In our implementation, this is solved using SSP, specifically the $k$-shortest-paths algorithm. As shown in [3], this algorithm has a worst-case time complexity of $\mathcal{O}(k(m+n\cdot\mathrm{log}\>n))$, where $k$ is the number of paths/tracks found by the algorithm and $m$ and $n$ refer to the number of nodes and edges, respectively, contained in the graph $\mathcal{G}^{\mathrm{trk}}$. The size and topology of $\mathcal{G}^{\mathrm{trk}}$ depends on $\mathcal{Y}$, and therefore is a function of both the size of the temporal window and the number of targets and false alarms.
In contrast, linear programming (LP)-based approaches to the min-cost flow problem have an average-case polynomial time complexity in the size of $\mathcal{G}^{\mathrm{trk}}$ [1]. This also applies to more recent approaches presented in [26] and [16], which embed the LP solution as a differentiable layer within a multi-layer perceptron. [^1] Note that [16] transforms the LP problem to a regularized quadratic program; however, the transformed problem has the same time complexity. The proposed SSP-GNN algorithm blends the edge-cost expressivity of learning-based methods with the reduced time complexity of SSP methods.

### v3

@@ I Introduction
Learning-based approaches to tracking [Xiang_learning_to_track_ICCV2015, Schulter_CVPR_2017] learn key components of a tracker (e.g., data association) directly from annotated training data. Recently, approaches to neural message passing [pmlr-v70-gilmer17a] have been adopted by [Braso_2020_CVPR] to formulate tracking on a detection graph (where a detection graph is instantiated for a temporal window given all the measurements/detections). However, predicting which pairs of detections should be linked without inferring complete tracks, while computationally efficient, may be suboptimal.
We propose an end-to-end bilevel formulation for learning MOT, where the inner optimization (tracking) is accomplished via a successive shortest paths (SSP) algorithm on a tracking graph [k_best_paths_castanon1990, Zhang_CVPR_2008] with edge costs. Given these costs, SSP finds a solution that is guaranteed to be optimal. In our outer optimization problem we learn parameters of a function that computes the cost of each edge in the tracking graph given detections that include attributes. The edge-cost prediction function takes the form of a graph neural network (GNN).
Learning GNN parameters requires solving a bilevel optimization problem that is different from those proposed earlier, e.g. [Schulter_CVPR_2017, li_learning_tracking_2022]. Specifically, we derive a novel loss function defined with respect to SSP-computed tracks and the ground-truth tracks and an algorithm to learn the GNN parameters by gradient descent. At each iteration, GNN parameters are updated to increase the cost of incorrect tracks, effectively learning from (tracking) mistakes.
@@
  • A novel end-to-end learnable approach to graph-based tracking with lower computational cost than [Schulter_CVPR_2017, li_learning_tracking_2022].
@@ II Related Work
Learning approaches to tracking In many domains, using machine-learning tools to find optimal parameters for a tracking algorithm from data have been shown to outperform hand-crafting and hand-tuning an algorithm. In the computer vision domain, learning-based approaches to multi-object tracking have taken many different forms. [Xiang_learning_to_track_ICCV2015] Demonstrated a Markov decision process (MDP) capable of initiating and terminating target tracks; parameters of this MDP were learned from data. In [rnn_tracking_milan2017] recurrent neural networks were used in an end-to-end framework. Specifically, in [rnn_tracking_milan2017] a target’s state was maintained by a dedicated recurrent neural network (RNN); assignment of measurements to target tracks and track creation was estimated via a long-short-term memory (LSTM) neural network. A transformer neural network was used in [Meinhardt_2022_CVPR] as a unified formulation for detecting (new) targets, extending tracks, and creating new tracks.
@@
Graph-based tracking via classical methods Application of graph-based and linear-programming (LP) methods to multi-target tracking has a long history, e.g., [k_best_paths_castanon1990], and includes recent attempts to combine graph-based tracking with multi-hypothesis tracking (MHT) [coraluppi_carthel_2019multiple]. Graph-based tracking has shown to be effective in computer vision where it is often employed with a tracking-by-detection paradigm where tracking is reduced to the grouping of detections (bounding boxes) proposed by a (target class specific) object detector.
@@
An early graph-based method of [Zhang_CVPR_2008] proposed the general form of a tracking graph and solved for an unknown number of targets via repeated invocation of min-cost flow. The computational complexity was reduced in [Pirsiavash_CVPR_2011] by introducing augmenting paths, thus extending an existing tracking solution to include an additional target instead of starting from scratch. While [Zhang_CVPR_2008] and [Pirsiavash_CVPR_2011] performed tracking in the image plane, tracking on the ground plane from multiple camera views via K-shortest-paths (KSP) algorithm was proposed in [berclaz_pami2011]. Approaches mentioned thus far assumed that targets dynamics are Markovian and targets move independently of each other. An approach to extend LP-based tracking to take into account long-term (group) behavior was proposed in [non_markovian_Maksai_ICCV_2017]. During training, their method alternated between learning behavior patterns and fitting trajectories to data given the observations and priors given by the learned patterns.
@@
Graph-based tracking via neural methods A significant limitation of early graph-based approaches [Zhang_CVPR_2008, Pirsiavash_CVPR_2011, berclaz_pami2011] is the requirement to hand-craft scalar costs on tracking graph edges. However, in many real-world scenarios where we wish to apply feature-aided tracking, measurements (detections) in addition to containing positional information also include (high-dimensional) vectors of attributes, such as re-identification (ReID) features. Neural networks could learn a mapping from these high-dimensional attributes to scalar edge weights, but deriving an end-to-end learnable method is nontrivial. Among the first attempts to apply neural methods to graph-based tracking, the approach of [Schulter_CVPR_2017] stands out for formulating learning and tracking as bilevel [intro_to_bilevel_opt_zhang2023] optimization. Specifically, in [Schulter_CVPR_2017] the high-level optimization problem solved for the neural network parameters, while the lower-level problem was a constrained LP that solved a tracking-by-detection problem over a temporal window of pre-defined length. The bilevel approach of [Schulter_CVPR_2017] was revisited in [li_learning_tracking_2022], but instead of relying on the implicit function theorem to deal with the inner LP problem, the latter used KKT conditions to compute the gradient at the lower problem’s optimum.
@@
Recent advances in neural methods for graph-structured problems, such as graph neural networks (GNNs) [defferard2016] led to advances in graph-based tracking. Inspired by the neural message-passing on graphs method of [pmlr-v70-gilmer17a], the authors of [Braso_2020_CVPR] proposed to learn a function to estimate edge probabilities in a tracking graph. This approach was extended in [Cetintas_braso_2023_CVPR] to jointly solve for short-term and long-term tracking. While the approaches of [Braso_2020_CVPR, Cetintas_braso_2023_CVPR] require a batch of frames, and thus operate with a constant lag, approaches of [Wang2020_GNNDetTrk, Meinhardt_2022_CVPR] require only the current frame to manage tracks. The ideas of [pmlr-v70-gilmer17a] were adopted in [liangnebp_fusion2022] to develop a neural-enhanced message-passing data association algorithm.
@@
Neural approaches to optimization Learning-based approaches to MOT mentioned thus far solve the underlying assignment problem using classical methods e.g., LP. An approach to solve quadratic optimization as a layer in a neural network was proposed in [optnet_pmlr-v70-amos17a]. In [sinkhorn_papakis2020gcnnmatch], the matrix of association costs was updated using Sinkhorn-Knopp’s algorithm [sinkhorn1967concerning] to satisfy MOT constraints. A differentiable graph matching layer was proposed in [learnable_graph_matching_He_CVPR2021] and used to associate graphs formed on detections across time frames.
@@
Our contribution We draw on [defferard2016, pmlr-v70-gilmer17a] and [Braso_2020_CVPR] to formulate a GNN that infers edge costs on a graph structure, but unlike [Braso_2020_CVPR] our inner optimization with SSP ensures that track predictions are globally optimal and satisfy constraints. We demonstrate in Section IV that our method compares favorably to a GNN-based method without such inner optimization. Our method is similar to that of [li_learning_tracking_2022] in forming a bilevel problem during training, but we show in Section III-E that our inner optimization with SSP has lower computational complexity.
@@ III-A Graph-based formulation
Within the batch of $T$ data frames, we solve the tracking problem by casting it in a graph-based formulation [Zhang_CVPR_2008, Pirsiavash_CVPR_2011]. Given a temporal window $t\in 1,\ldots,T$ our objective is to group $\{{\bf y}_{j,t}\}$ into a set $\mathcal{P}\doteq\{p_{k}\}$ of paths in a tracking graph $\mathcal{G}^{\textrm{trk}}$ with positive and negative edge costs. The set $\mathcal{P}$ may be empty if the tracker doesn’t predict any tracks.
@@ III-B The detection graph and the tracking graph
Given detections $\mathcal{Y}\doteq\{{\bf y}_{j,t}\}_{t=1}^{T}$ in a temporal window with $T$ frames, we construct a detection graph $\mathcal{G}^{\textrm{det}}$. Nodes in $\mathcal{G}^{\textrm{det}}$ are in one-to-one correspondence with detections in $\mathcal{Y}$. An edge $(j,j^{\prime})$ exists between ${\bf y}_{j,t}$ and ${\bf y}_{j^{\prime},t^{\prime}}$ if $t<t^{\prime}\leq t+\Delta$ and the location components of ${\bf y}_{j,t}$ and ${\bf y}_{j^{\prime},t^{\prime}}$ satisfy gating constraints. The use of $\Delta>1$ is standard [Schulter_CVPR_2017, li_learning_tracking_2022] to handle short-duration obscuration (occlusion) of a target. In Fig. 2(a) we show an example of $\mathcal{G}^{\textrm{det}}$ for $T=3$ and $\Delta=2$. Two paths that satisfy node-disjoint constraints are shown.
Given $\mathcal{G}^{\textrm{det}}$, we define its companion tracking graph $\mathcal{G}^{\textrm{trk}}$, whose topology follows [k_best_paths_castanon1990, Zhang_CVPR_2008]. In particular, every node in $\mathcal{G}^{\textrm{det}}$ is represented by a pair of “twin” nodes in $\mathcal{G}^{\textrm{trk}}$. Furthermore, $\mathcal{G}^{\textrm{trk}}$ contains source $\bf s$} and terminal $\bf t$} nodes with edges from $\bf s$} and to $\bf t$} defined in a way as to allow every detection to either correspond to a tracks’s birth or its termination. Fig. 2(b) shows the tracking graph $\mathcal{G}^{\textrm{trk}}$ derived from its $\mathcal{G}^{\textrm{det}}$; the paths (tracks) in $\mathcal{G}^{\textrm{trk}}$ originate in $\bf s$}, pass through “twin” nodes and terminate in $\bf t$}.
@@
Our $f_{\boldsymbol{\theta}}$ takes the form of a message-passing network (MPN) [pmlr-v70-gilmer17a], a specialization of a graph neural net (GNN) [defferard2016]. We adopt the ideas of [pmlr-v70-gilmer17a] and  [pmlr-v80-kipf18a] as implemented in [Braso_2020_CVPR].
@@
1Initialize ${\bf h}^{(0)}_{(i,j)}$ and ${\bf h}^{(0)}_{z}$ for all $(i,j)\in\mathcal{E},z\in\mathcal{V}$
@@ 18 $c_{(i,j)}=g_{e}^{\mathrm{readout}}({\bf h}^{(L)}_{(i,j)},{\bf h}^{(L)}_{(j,i)})$
Messages are computed using the functions $g_{e}(\cdot,\cdot,\cdot)$ and $g_{v}(\cdot,\cdot)$ which take the form of learnable multi-layer perceptions. In our implementation, as suggested by [Braso_2020_CVPR], $g_{v}(\cdot,\cdot)$ consists of two independently parameterized functions $g_{v}^{\mathrm{past}}(\cdot,\cdot),g_{v}^{\mathrm{fut}}(\cdot,\cdot)$ which are specialized for forward and backward in time messages respectively. The readout function $g_{e}^{\mathrm{readout}}(\cdot,\cdot)$, also a learnable MLP, then computes the edge cost by aggregating the final attributes of the forward and backward in time edges between each pair of nodes. The parameters of these four functions comprise the vector $\boldsymbol{\theta}$}.
@@ III-C End-to-end learning for tracking
Bilevel optimization to learn MPN parameters We formulate the following bilevel optimization problem [colson_overview_bilevel_2007, intro_to_bilevel_opt_zhang2023, bilevel_opt_nonsmooth_ochs2015] to find the optimal set of GNN parameters given ground-truth paths $\mathcal{P}^{+}$ :
@@
$\dfrac{\partial\mathcal{L}}{\partial{\boldsymbol{\theta}}}=\dfrac{\partial\mathcal{L}}{\partial{\bf c}}\Big|_{\mathcal{P}^{*}}\dfrac{\partial{\bf c}}{\partial{\boldsymbol{\theta}}}\;,$ (4)
@@
In practice, we implement optimization of $\boldsymbol{\theta}$} in Equation 3 via a two-stage learning algorithm. In Stage I, inspired by [oh_sastry_mcmcda_2009] which introduced MCMC data association moves, we “perturb” $\mathcal{P}^{+}$ to pre-compute a fixed-size set of negative examples $\mathcal{P}^{-}$. Generating $\mathcal{P}^{-}$ does not require running the SSP algorithm.
@@ 9 ${\boldsymbol{\theta}}\leftarrow\mathsf{init\_gnn\_weights}()$
11
12  0.2em $\{c_{e}\}_{e=1}^{|\mathcal{E}|}\leftarrow f_{\boldsymbol{\theta}}(\mathcal{G})$
@@
18
19  0.2em $\{c_{e}\}_{e=1}^{|\mathcal{E}|}\leftarrow f_{\boldsymbol{\theta}}(\mathcal{G})$
@@
III-D Tracking with a trained model
Once training is complete, the model weights are frozen and inference is performed without a backward pass, i.e. the bottom arrow of Fig. 1. Our experiments evaluate model performance over individual tracking windows. Our method can be extended to long tracking sequences by splitting them into smaller temporal windows and applying standard methods for re-initializing and stitching tracks across overlapping windows [Braso_2020_CVPR], [li_learning_tracking_2022].
III-E Computational Complexity
In our analysis, to provide a fair comparison between our proposed algorithm and related methods we focus only on forward inference of the inner optimization problem used to compute $\mathcal{P}^{\star}$ from $\mathcal{G}^{\mathrm{trk}}$. In our implementation, this is solved using SSP, specifically the $k$-shortest-paths algorithm. As shown in [berclaz_pami2011], this algorithm has a worst-case time complexity of $\mathcal{O}(k(m+n\cdot\mathrm{log}\>n))$, where $k$ is the number of paths/tracks found by the algorithm and $m$ and $n$ refer to the number of nodes and edges, respectively, contained in the graph $\mathcal{G}^{\mathrm{trk}}$. The size and topology of $\mathcal{G}^{\mathrm{trk}}$ depends on $\mathcal{Y}$, and therefore is a function of both the size of the temporal window and the number of targets and false alarms.
In contrast, linear programming (LP)-based approaches to the min-cost flow problem have an average-case polynomial time complexity in the size of $\mathcal{G}^{\mathrm{trk}}$ [Ahuja1993NetworkFT]. This also applies to more recent approaches presented in [Schulter_CVPR_2017] and [li_learning_tracking_2022], which embed the LP solution as a differentiable layer within a multi-layer perceptron. [^1] Note that [li_learning_tracking_2022] transforms the LP problem to a regularized quadratic program; however, the transformed problem has the same time complexity. The proposed SSP-GNN algorithm blends the edge-cost expressivity of learning-based methods with the reduced time complexity of SSP methods.